# TEncDM: Understanding the Properties of Diffusion Model in the Space of Language Model Encodings

### v1

@@ 9 Limitations
There are two limitations that warrant further investigation. First, while the quality of the model can be improved by training diffusion encoder, decoder and denoising model simultaneously, we avoid doing so in order to avoid overcomplicating the approach. Second, samples from the latent space have a high dimensionality that depends on the sequence length, making the training of our method significantly slower as the length increases. This problem can probably be eliminated by training the autoencoder, which is a great direction for the further research.
@@ 10 Conclusion
In this work, we explore key details of the diffusion pipeline for text generation. We propose TEncDM which trains the diffusion model inside the latent space of language encoder model. In order to improve text generation performance, we analyse the effect of self-conditioning and conclude that it increases the magnitudes of model’s predictions and results in reducing of generation steps. We also propose an efficient decoder that boosts the diffusion model performance. The extensive ablation on ROCStories proves the impact of proposed design choices. TEncDM outperforms recent diffusion models, non-autoregressive and classical autoregressive methods thorough experiments on downstream tasks.
@@ Appendix A Decoder for embedding-based model
We show that our proposed decoder is robust not only for encoding-based diffusion model, but also for embedding-based one. In Table 6, we compare our decoder described in Section 7.2 with the commonly used rounding to the closest embedding. It is easy to see that our decoder improves the text quality according to MAUVE. Also, it hugely improves Memorization and Diversity. Low value of Perplexity for the rounding method comes from the low diversity and it does not imply the high quality of the generated samples.
@@ Appendix B Corruption for decoder training
The both techniques introduce the random noise into the decoder input. However, the first one attempt to mimic the samples from the diffusion model denoising trajectory. We implement it by randomly sampling the timestep from the range $t\in[0,t_{m}ax]$ and calculating the corresponding $z_{t}$. In Figure 6, we show the text generation quality in terms of Perplexity and MAUVE Score with respect to $t_{m}ax$. In Figure 5, we present the similar result for the second decoder training technique with varying noise strength $\sigma$. To make the comparison fair we apply all decoders to the same latents produced by the diffusion model. Both plots suggest that there is an optimal amount of noise that should be added. However, the first technique results in a better performance.
@@
Appendix C Self-conditioning increases prediction magnitude
We show that self-conditioning tend to increase the magnitude of values of model’s output by conducting the following experiment. We sample $z_{t}$ using the diffusion forward process and predict $\tilde{z}_{0}^{t}=\hat{z}_{\theta}(z_{t},t,\tilde{z}_{0}^{t})$ from it several times. Each time we feed the model its previous prediction and do not change $z_{t}$ and timestep $t$. In Figure 7, we plot the trajectories of prediction magnitude obtained by this repeated prediction scheme for different timesteps $t$. The results show that the prediction magnitude grows at each step, even though we change only the sample, which we provide to a model using the self-conditioning. This allows us to conclude that self-conditioning is indeed responsible for the increase in prediction magnitude, which is reflected in the inference behaviour of the model.
@@ Appendix E Implementation details
We train our models using 4 A100 GPUs. The training takes approximately 10 hours for ROCStories, 10 hours for QQP and 30 hours for XSum.
ROCStories XSum QQP
@@
Batch Size 512
Warmup Steps 500
@@
Training Steps 100k 50k 100k
Max Seq Length 80 64 64
Max Context Length 256 32
@@ Appendix F Dataset Statistics
The dataset is used for summarization task and it contains 204k BBC articles, which are provided as document and summary pairs and covered wide range of topics (Sports, Politics, etc.). It has 204,045 training instances, 11,332 validation instances, and 11,334 test instances.
@@
The subset of QQP dataset, proposed in Gong et al. (2023), consists of 144k question pairs from the Quora platform that are paraphrases of each other. It has 144,715 training instances, 2,048 validation instances, and 2,500 test instances.
Appendix G Comparison with GPT2
We compare our diffusion model with fine-tuned GPT2-small Radford et al. (2019) on an unconditional generation task using ROCStories Mostafazadeh et al. (2016) dataset. We use the Nucleus sampling with $p=0.9$ for the GPT generation, as is produced the best results. Both models have similar amount of parameters (124M for GPT2 and 145M for TEncDM). The result of the comparison is presented in Table 8 and it shows that GPT2 has a higher MAUVE, but it also tends to memorise the training data set more and has a lower diversity. The perplexity comparison is unfair as it is computed with the GPT2-large model, which behaves similarly to GPT2-small. Given that the GPT2 is pre-trained and TEncDM was trained from scratch, we can conclude that both models perform at about the same level.
Decoder ppl $\downarrow$ mem $\downarrow$ div $\uparrow$ mauve $\uparrow$
GPT2-small FT $15.5_{.11}$$519_{.004}.$$269_{.003}.$$739_{.031}.$
TEncDM $34.1_{.66}$$412_{.005}.$$304_{.006}.$$707_{.024}.$
Table 8: Comparison on unconditional generation (ROCStories).
Appendix H Generation examples
BERT enc
with
MLP
decoder
##ocks
A man wanted to go swimming. They packed up the boat, and drove to the beach. They found a nice spot by the water. They swam for hours, remorving the scenery. At the end of the trip, they had to go home. chantingctic Widow leopard paranoidntialivatingolar chanting Xiaocticntlymurananteurboopectatorctictleshalanadonantnantnadonantavesmura
##mps
Heather wanted to bake a cake. She grabbed some ingredients and put the cake in the oven. Her alarm rang, but didn’t go off. To her dismay the cake was on fire! The cake was so mess that she forgot to turn off the oven.putmpsmuravatednantpectatoraves Wan emitted chantingmura leopardmura leopardave sputvatednantavesnant Widownantnantavesshing
##eur
Rita was always bullied in school. But every time time she stood up, she was bullied. Rita was too young. But as the bully grew, she improved. After school, Rita was no longer bullied. fraction Signlatingbeknantbekaveslaxivatingmpsivatingmpspectatoromoomonadoavespectatoravessh ingavesmpsolarivatingutavatedivating Widowoveavesriotmpsmps
BERT emb
with
Transformer
decoder
@@ Appendix H Generation examples
Last week my brother brought my skateboard with me. He started using the skateboard after half an hour long. I bit my leg and started to fall out of my foot. My brother got into the piece. He was able to scolded me and take me to the hospital.
@@
BERT enc
with
Transformer
decoder
Emily wanted her nails become pink. She took some nailolish from a grocery store and thought it looked horrible. She tried everything to get rid of it. It ended up making a ton of mess. Emily had to throw the mess all out.
@@
T5 enc
with
Transformer
decoder
Jack had a dog that he loved named Frankt. He was a big Shepherd who had lotss barkles and collar. One day, Jack left Fredt at his house and didn t find him. After three days, Ft’s owner found out. bought a searches. The next day, his owner found Frankt in the house.
The kids climbed outside with the gun. They wanted to shooting their neighbor’ to gun. They fell on a higher of a mountain. mom tried to carry the rifle for them. It was too heavy to carry from the kids.
Shera and her weddingrs were packing a box of pictures. Shera du searched through each box for the favorite picture. Finally it was time stamped the numbers. Shera put the pictures in the box in front of the machine..a is, it took of lot time to up the number right.
Table 9: Examples of generated texts for different models.

### v2

@@ 9 Limitations
There are three limitations that warrant further investigation. First, while the quality of the model can be improved by training diffusion encoder, decoder and denoising model simultaneously, we avoid doing so in order to avoid overcomplicating the approach. Second, samples from the latent space have a high dimensionality that depends on the sequence length, making the training of our method significantly slower as the length increases. This problem can probably be eliminated by training the autoencoder, which is a great direction for the further research. Third, as different encoders works better for different tasks, it is necessary to find the best one for each task.
@@ 10 Conclusion
In this work, we explore key details of the diffusion pipeline for text generation. We propose TEncDM which trains the diffusion model in the latent space of language model encoding. In contrast to embeddings, they contain contextual information which helps diffusion model to recover latents. In order to improve text generation performance, we analyse the effect of self-conditioning and conclude that it increases the magnitudes of the model’s predictions and results in reducing of generation steps number. Additionally, we propose an efficient decoder and noise scheduler that boost the diffusion model performance. The extensive ablation on ROCStories and Wikipedia proves the impact of proposed design choices. Finally, TEncDM outperforms recent diffusion models and classical autoregressive methods in downstream task experiments.
@@ Appendix A Decoder for embedding-based model
We show that our proposed decoder is robust not only for encoding-based diffusion model, but also for embedding-based one. In Table 5, we compare our decoder described in Section 7.2 with the commonly used rounding to the closest embedding on ROCStories dataset. It is easy to see that our decoder improves the text quality according to mauve. Also, it hugely improves memorization and diversity. Low value of perplexity for the rounding method comes from the low diversity and it does not imply the high quality of the generated samples.
@@ Appendix B Corruption for decoder training
Both techniques introduce a random noise into the decoder input. However, the first one attempts to mimic samples from the diffusion model denoising trajectory. We implement it by randomly sampling a timestep from the range $t\in[0,t_{max}]$ and calculating the corresponding $z_{t}$. In Figure 5, we show the text generation quality in terms of Perplexity and MAUVE Score with respect to $t_{max}$. In Figure 6, we present the similar result for the second decoder training technique with varying noise strength $\sigma$. To make the comparison fair we apply all decoders to the same latents produced by the diffusion model. Both plots suggest that there is an optimal amount of noise that should be added. However, the first technique results in a better performance.
@@
Appendix C Self-conditioning
C.1 Self-conditioning increases prediction magnitude
We show that self-conditioning tend to increase the magnitude of values of model’s output by conducting the following experiment on ROCStories dataset. We sample $z_{t}$ using the diffusion forward process and predict $\tilde{z}_{0}^{t}=\hat{z}_{\theta}(z_{t},t,\tilde{z}_{0}^{t})$ from it several times. Each time we feed the model its previous prediction and do not change $z_{t}$ and timestep $t$. In Figure 7, we plot the trajectories of prediction magnitude obtained by this repeated prediction scheme for different timesteps $t$. The results show that the prediction magnitude grows at each step, even though we change only the sample, which we provide to a model using the self-conditioning. This allows us to conclude that self-conditioning is indeed responsible for the increase in prediction magnitude, which is reflected in the inference behaviour of the model.
@@ C.1 Self-conditioning increases prediction magnitude
C.2 Wikipedia self-conditioning analysis
We conduct the experiments from Section 7.2 on Wikipedia dataset to demonstrate the generalizability of self-conditioning effect. Figure 8 illustrates that with self-conditioning the MAUVE performance initially increases with the number of generation steps, reaching a certain level. Thereafter, it begins to decline. This behavior is similar to that observed on the ROCStories dataset. In contrast, for the model without self-conditioning, the quality exhibits a monotonic increase. However, unlike the model trained on ROCStories, the Wikipedia-trained model achieves similar performance for both 50 and 100 generation steps.
Figure 8: Comparison of models with and without self-conditioning on Wikipedia dataset.
Figure 9 depicts the trajectories of prediction magnitudes for varying amounts of generation steps in the Wikipedia dataset. It can be observed that the trajectories exhibit a similar trend to those observed in ROCStories, although the spread of final prediction magnitudes is larger.
Figure 9: Comparison of prediction magnitudes for generation processes with different amount of steps on Wikipedia dataset.
@@ Appendix D Noise Schedulers
In Figure 10, we visualize the noise addition rate in the forward process in terms of $\sqrt{\alpha_{t}}$ for different noise schedulers.
@@ Appendix E Implementation details
We train our models using 4 A100 GPUs. The training takes approximately 6 hours for ROCStories, 15 hours for Wikipedia, 1 hour for QQP, 6 hours for XSum and 6 hours for Wiki-Auto. Table 6 shows hyperparameters for each dataset. For each dataset we chose the largest learning rate which keeps the training stable and train the model until convergence. The batch size for XSum was chosen smaller to fit on the GPU. We do not tune other hyperparameters, even though it might improve quality.
ROCStories Wikipedia XSum QQP Wiki-Auto
@@
Decoder Trainable Params 44M
Decoder Self-Attention Heads 16
Decoder Transformer Layers 3
@@
Warmup Steps 5000
@@
Batch Size 512 512 128 512 512
Training Steps 200k 500k 100k 200k 200k
Max Seq Length 80 128 64 50 64
Max Context Length 512 50 64
Sampling steps 50
Guidance coefficient 0.5 0.5 0.5
@@ Appendix F Dataset Statistics
Wikipedia
For large-scale experiments, we utilize the English Wikipedia subset from the ROOTS corpus (Laurençon et al. 2023). Additionally articles containing fewer than 600 characters are removed, and the remaining articles are segmented into sequences of approximately 128 tokens each. The dataset comprises a total of 15.4M sequences, of which 10,000 are allocated as a validation set.
@@
This dataset is used for summarization task and it contains 204k BBC articles, which are provided as document and summary pairs and covered wide range of topics (Sports, Politics, etc.). It has 204,045 training instances, 11,332 validation instances, and 11,334 test instances.
@@
The subset of QQP dataset, proposed in (Gong et al. 2023), consists of 144k question pairs from the Quora platform that are paraphrases of each other. It has 144,715 training instances, 2,048 validation instances, and 2,500 test instances.
Wiki-Auto
The dataset proposed by (Jiang et al. 2020) is designed for solving the text simplification task. The goal is to turn complex text into sequences with simplified grammatical structures and a more limited range of vocabulary. We use the preprocessing scheme outlined by (Gong et al. 2023), resulting in a dataset with 677,000 training instances, 2,050 validation instances, and 5,000 test instances.
Appendix G Low MAUVE for RoBERTa encoder
Table 1 illustrates that, despite satisfactory perplexity, memorialization, and diversity values, the MAUVE metric exhibits a low value for RoBERTa encoder. This outcome is not attributable to inferior text quality but largely to the discrepancy in the lengths of the generated and validation texts. MAUVE is estimated based on GPT2 encoding of the text, which extracts information about text length, among other elements. To validate this hypothesis, two experiments were conducted.
Experiment 1
In Figure 11, we visualize the text length distributions for TEncDM with RoBERTa and BERT encoders, and for the source text. It is clear that the BERT’s distribution is much more similar to the source text distribution than RoBERTa’s. This suggests that the reason for drop in MAUVE value might indeed be the text length.
Figure 11: Text length distribution for source texts and TEncDM generation with RoBERTa and BERT encoders.
Experiment 2
To match the length distributions, we truncate all generated and source texts to 40 tokens and measure MAUVE for them. The measurement results are summarized in Table 7. The obtained values increased significantly compared to the values in Table 1, and also became closer to each other. This indicates that MAUVE is indeed dependent on the length of texts.
Encoder mauve $\uparrow$
BERT .$940_{.013}.$
RoBERTa .$913_{.009}.$
Table 7: MAUVE for TEncDM with BERT and RoBERTa encoders on texts truncated to 40 tokens.
Experiment 3
According to the MAUVE calculation algorithm, we first cluster vector representations of texts using K-Means. Then for both generated and source texts we construct the histogram of cluster sizes based on the resulting clustering. If the histograms are similar, it can be inferred that the generated and source texts have similar distribution, which indicates a high degree of generation quality. Conversely, if the histograms are different, it can be concluded that the distribution of generated texts is distinct from the distribution of source texts. Such histograms for RoBERTa encoder are shown in Figure 12.
Figure 12: Histograms used for MAUVE calculation.
We can analyze the differences in the distributions by comparing clusters, wherein the number of generated and real texts is balanced, with unbalanced clusters.
First, we calculate the entropy of the distribution of generated and real texts in each cluster to measure its balance. The resulting entropies are then sorted to yield a ranked list of clusters. Then, as we hypothesise that the main difference between texts is their length, we calculate the average text length in each cluster in terms of tokens. We take in consideration only those clusters, in which there are more generated texts than source ones. We plot these lengths sorted by cluster entropy in Figure 13 for both RoBERTa and BERT encoders. We have applied smoothing to each plot in order to enhance the visibility of the trend.
Figure 13: Smoothed average text lengths for clusters sorted by entropy.
In Figure 13, the entropy increases from left to right with the cluster rank, which means that the clusters on the left are more imbalanced than those on the right. We observe that the length of RoBERTa’s texts decreases with an increase in balance, whereas the length of BERT’s texts remains relatively constant. This suggests that RoBERTa clusters without source texts tend to have longer texts, which is the underlying reason for the decline in MAUVE performance.
Appendix H Scalability
We investigate the scalability of TEncDM. Our model leverages both a text encoder and a diffusion model, offering two avenues for parameter expansion: scaling the diffusion model alone or scaling both the encoder and diffusion model concurrently. In this section, we focus on the first approach, where we exclusively increase the size of the diffusion model. Specifically, we augmented the number of transformer heads from 12 to 16 and doubled the number of layers from 12 to 24. We call the standard model “base” and the increased model “large”. In Table 8 we show comparison of base and large diffusion models on Wikipedia dataset. The results demonstrate the effectiveness of the scaling strategy. The second approach, involving simultaneous scaling of both the encoder and diffusion model, necessitates significantly greater computational resources and training time. Therefore, we defer the exploration of this approach to future research.
Model size ppl $\downarrow$ mem $\downarrow$ div $\uparrow$ mauve $\uparrow$
large $84.5_{0.8}$$298_{.001}.$$505_{.002}.$$909_{.011}.$
base $104.4_{2.1}$$286_{.002}.$$504_{.003}.$$874_{.011}.$
Table 8: Scalability of the diffusion model on Wikipedia.
@@ Appendix H Scalability
BERT embeddings with Transformer decoder    Last week my brother brought my skateboard with me. He started using the skateboard after half an hour long. I bit my leg and started to fall out of my foot. My brother got into the piece. He was able to scolded me and take me to the hospital.
@@
BERT encodings with Transformer decoder    Emily wanted her nails become pink. She took some nailolish from a grocery store and thought it looked horrible. She tried everything to get rid of it. It ended up making a ton of mess. Emily had to throw the mess all out.
@@
Table 9: Examples of generated texts for different models on ROCStories dataset. Generation inaccuracies are highlighted.
Appendix I Comparison with GPT-2
We compare TEncDM with fine-tuned GPT-2-small (Radford et al. 2019) on an unconditional generation task using ROCStories and Wikipedia datasets. We use the Nucleus sampling with $p=0.95$ for the GPT-2 generation, as it produced the best results. Both models have similar amount of parameters (124M for GPT-2 and 145M for TEncDM). The result of the comparison is presented in Table 10 and it shows that our approach suprasses fine-tuned GPT-2-small by all metrics except perplexity, because GPT-2 tends to memorise the training data set more and has a lower diversity. However, the perplexity comparison is unfair as it is computed with the GPT-2-large model, which behaves similarly to GPT-2-small.
Decoder ppl $\downarrow$ mem $\downarrow$ div $\uparrow$ mauve $\uparrow$
ROCStories
GPT2-small FT $\textbf{15.5}_{.11}$$519_{.004}.$$269_{.003}.$$739_{.031}.$
TEncDM $29.1_{.89}$${\textbf{.453}}_{.003}.$${\textbf{.295}}_{.002}.$$\textbf{.762}_{.043}.$
Wikipedia
GPT2-small FT $\textbf{15.3}_{.18}$$508_{.004}.$$380_{.003}.$$708_{.025}.$
TEncDM $104.4_{2.1}$$\textbf{.286}_{.002}.$$\textbf{.504}_{.003}.$$\textbf{.874}_{.011}.$
Table 10: Comparison on unconditional generation.
Appendix J Generation examples
We show generation examples for diffusion model trained in the latent space of embeddings and encodings in Table 9.